# Epidemic threshold and localization of the SIS model on directed complex networks

### v1

@@ IV.1 Linear stability analysis
For $c>c_{\rm gap}$, $\Lambda_{1}=c\lambda$ is an outlier Neri and Metz (2016), and the trivial solution is stable if $c<\lambda^{-1}$. For $c\leq c_{\rm gap}$, the spectral gap is zero, $\Lambda_{1}\in\mathbb{R}$ belongs to the boundary of the continuous spectrum, and the trivial fixed-point is stable for
@@
These results for the stability analysis are universal, as they are independent of higher moments of the distributions $p_{k\ell}$ and $P_{\lambda}(x)$ characterizing the network structure.
@@ IV.2 Phase diagram
Figure 1: Comparison between the solutions of Eq. (11) (solid lines) with the fixed-point solutions of Eq. (8) (symbols) for directed networks with a $\Gamma$-distribution of infection rates and a Poisson degree distribution. The fixed-point solutions are derived from an ensemble of 5050 networks with $N=10^{5}$ nodes, while the population dynamics results are obtained from 55 independent runs of the algorithm with $M=10^{6}$ stochastic variables. (a) The prevalence $\langle\rho\rangle$ as a function of the mean infection rate $\lambda$ for standard deviation $\sigma=0.2$ of the infection rates. (b) and (c): distribution $\mathcal{P}(\rho)$ of the infection probabilities for $\lambda=1/2$, $\sigma=0.2$, and two different $c$.
@@
Figure 2: Phase diagram of the SIS model on directed networks in terms of the mean degree $c$ and the standard deviation $\sigma$ of the infection rates (the mean infection rate is $\lambda=1/2$). The indegrees follow a Poisson distribution, while the infection rates follow a $\Gamma$-distribution. The model exhibits an endemic phase ($\langle\rho\rangle>0$) and an absorbing phase ($\langle\rho\rangle=0$). The standard deviations at the dots are, respectively, $\sigma_{*}=\sqrt{\lambda(1-\lambda)}$ and $\sigma_{4}\simeq 0.36$. For $\sigma>\sigma_{*}$, the critical line is obtained by solving Eq. (11) using the population dynamics algorithm with $M=10^{6}$ stochastic variables. The colour scale shows the inverse participation, Eq. (38), which quantifies the spatial fluctuations of the infection probabilities. The fourth moment $\mu_{4}$ of the infection rates is defined in Eq. (42).
Figure 3: Prevalence $\langle\rho\rangle$ as a function of the mean degree $c$ for different distributions of infection rates and indegrees in the regime $\sigma<\sigma_{*}$ (see the main text). The infection rates have mean $\lambda=1/2$ and standard deviation $\sigma=0.2$. The results are obtained by solving Eq. (11) using the population dynamics algorithm with $M=10^{5}$ stochastic variables. The inset shows the prevalence near $c=\lambda^{-1}$ in logarithmic scale. The colours in the inset correspond to the same distributions as in the main panel.
@@ IV.3 Localization of epidemic spreading
Figure 5: Dimensionless second moment $\mathcal{I}^{(2)}(\rho)$ [Eq. (38)] and inverse participation ratio $\mathcal{I}^{(4)}(\rho)$ [Eq. (39)] as functions of the standard deviation $\sigma$ of the infection rates for different mean degrees $c$ close to the epidemic threshold $c=\lambda^{-1}$. These results are for directed networks with Poisson indegrees and a $\Gamma$-distribution of infection rates with mean $\lambda=1/2$. Symbols represent numerical results obtained from Eq. (11) using the population dynamics algorithm with $M=10^{6}$ (vertical bars indicate the standard deviation of the mean computed over 1010 independent runs). Solid lines are analytic predictions derived from the moments of the leading eigenvector [Eqs. (40) and (41)].
@@
of the infection rates. Depending on the network parameters, Eqs. (40) and (41) diverge due to vanishing denominators Metz and Neri (2021). In particular, $\mathcal{I}^{(4)}(R_{1})$ diverges at $c=\mu_{4}^{\frac{1}{3}}$. This naturally raises the question of whether $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$ exhibit a similar behaviour near $c=\lambda^{-1}$.
In figure 5, we show $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$ as functions of $\sigma$ near the epidemic threshold. As a comparison, this figure also displays Eqs. (40) and (41) at $c=\lambda^{-1}$, demonstrating that $\mathcal{I}^{(2)}(R_{1})$ and $\mathcal{I}^{(4)}(R_{1})$ diverge at $\sigma_{*}=\sqrt{\lambda(1-\lambda)}$ and $\sigma_{4}\simeq 0.36$, respectively, for the $\Gamma$-distribution of infection rates. The results for $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$ are overall consistent with the analytic predictions obtained from the moments of $\bm{R}_{1}$. The discrepancies observed in figure 5 stem from the slow convergence of the population dynamics algorithm near the thresholds $\sigma_{*}$ and $\sigma_{4}$, which makes it difficult to accurately determine the distribution $\mathcal{P}(\rho)$ from the solutions of Eq. (11). In addition, we note that $\mathcal{I}^{(4)}(\rho)$ remains bounded for $\sigma_{4}<\sigma<\sigma_{*}$, a consequence of the finite number of stochastic variables used to discretize $\mathcal{P}(\rho)$ in the population dynamics method.
@@ V Final remarks
In this work, we have determined the phase diagram of the SIS model on directed complex networks within the quenched mean-field approximation. By combining random-matrix results with an analytic approach for the distribution of stationary infection probabilities, we have computed the epidemic threshold as a function of the mean degree $c$ and the standard deviation $\sigma$ of the infection rates defining the contact network. Our results show that the SIS model exhibits a transition between the absorbing and endemic phases provided $c\geq\lambda^{-1}$, where $\lambda$ is the average infection rate.
Remarkably, the critical line is independent of the degree distribution but it depends strongly on the distribution of infection rates. When $\sigma<\sigma_{*}=\sqrt{\lambda(1-\lambda)}$, both the epidemic threshold and the critical behaviour are governed by the leading eigenpair of the contact network Metz and Neri (2021) and are independent of the infection-rate distribution. In contrast, for $\sigma>\sigma_{*}$, the fluctuations of infection rates have a pronounced effect on the critical line. While the SIS model undergoes an absorbing phase transition for a $\Gamma$-distribution of infection rates as $\sigma$ increases, it remains in the endemic phase for a Pareto distribution. This striking difference is explained by the large fraction of near-zero infection rates generated by the $\Gamma$-distribution, which suppresses the endemic state for large $\sigma$. Together, these results provide a comprehensive picture of how network heterogeneity shapes the phase diagram of the SIS model on directed networks.
@@
We remark that our results are strictly valid in the limit $N\rightarrow\infty$, as they are primarily based on the numerical solutions of Eq. (11). Consequently, our approach does not address how the infection probabilities or the number of potentially infected nodes scale with the system size, which could in principle clarify, for instance, why the dimensionless second moment remains finite as the IPR diverges. A detailed analysis of this issue, and its relation to the finite-size scaling of the infection probabilities, is an interesting direction for future work.
From a methodological perspective, we have introduced a general analytic framework to obtain an equation for the distribution of fixed-point states in the limit $N\rightarrow\infty$. The method is not restricted to the SIS model and can be extended to study the non-equilibrium fixed-points of other dynamical systems on directed networks, including the Kuramoto model Acebrón et al. (2005) and firing-rate models of neural networks Sompolinsky et al. (1988). Moreover, since the approach is based on the cavity method, it can also incorporate other important network features, such as short loops and degree-degree correlations, opening promising avenues of future research.

### v2

@@ IV.1 Linear stability analysis
For $c>c_{\rm gap}$, $\Lambda_{1}=c\lambda$ is an outlier Neri and Metz (2016), and the trivial solution is stable if $c<\lambda^{-1}$. For $c\leq c_{\rm gap}$, the spectral gap is zero, $\Lambda_{1}=\sqrt{c\left(\sigma^{2}+\lambda^{2}\right)}\in\mathbb{R}$ belongs to the boundary of the continuous spectrum, and the trivial fixed-point is stable for
@@
These results for the stability analysis hold for $c>1$. The stability thresholds are universal, as they are independent of higher moments of the distributions $p_{k\ell}$ and $P_{\lambda}(x)$ characterizing the network structure.
@@ IV.2 Phase diagram
Figure 1: Comparison between the solutions of Eq. (11) (solid lines) with the fixed-point solutions of Eq. (8) (symbols) for directed networks with a Poisson degree distribution and a $\Gamma$-distribution of infection rates or coupling strengths. The fixed-point solutions are derived from an ensemble of 5050 networks with $N=10^{5}$ nodes, while the population dynamics results are obtained from 55 independent runs of the algorithm with $M=10^{6}$ stochastic variables. (a) The prevalence $\langle\rho\rangle$ as a function of the mean infection rate $\lambda$ for standard deviation $\sigma=0.2$ of the infection rates. (b) and (c): distribution $\mathcal{P}(\rho)$ of the infection probabilities for $\lambda=1/2$, $\sigma=0.2$, and two different $c$.
@@
Figure 2: Phase diagram of the SIS model on directed networks in terms of the mean degree $c$ and the standard deviation $\sigma$ of the infection rates or coupling strengths (the mean infection rate is $\lambda=1/2$). The indegrees follow a Poisson distribution, while the infection rates follow a $\Gamma$-distribution. The model exhibits an endemic phase ($\langle\rho\rangle>0$) and an absorbing phase ($\langle\rho\rangle=0$). The standard deviations at the dots are, respectively, $\sigma_{*}=\sqrt{\lambda(1-\lambda)}$ and $\sigma_{4}\simeq 0.36$. For $\sigma>\sigma_{*}$, the critical line is obtained by solving Eq. (11) using the population dynamics algorithm with $M=10^{6}$ stochastic variables. The colour scale shows the inverse participation, Eq. (45), which quantifies the spatial fluctuations of the infection probabilities. The fourth moment $\mu_{4}$ of the infection rates is defined in Eq. (49).
Figure 3: Prevalence $\langle\rho\rangle$ as a function of the mean degree $c$ for different distributions of infection rates and indegrees in the regime $\sigma<\sigma_{*}$ (see the main text). The infection rates or coupling strengths have mean $\lambda=1/2$ and standard deviation $\sigma=0.2$. The results are obtained by solving Eq. (11) using the population dynamics algorithm with $M=10^{5}$ stochastic variables. The inset shows the prevalence near $c=\lambda^{-1}$ in logarithmic scale. The colours in the inset correspond to the same distributions as in the main panel.
@@ IV.3 Localization of epidemic spreading
Figure 5: Dimensionless second moment $\mathcal{I}^{(2)}(\rho)$ [Eq. (46)] and inverse participation ratio $\mathcal{I}^{(4)}(\rho)$ [Eq. (45)] as functions of the standard deviation $\sigma$ of the infection rates for different mean degrees $c$ close to the epidemic threshold $c=\lambda^{-1}$. These results are for directed networks with Poisson indegrees and a $\Gamma$-distribution of infection rates or coupling strengths with mean $\lambda=1/2$. Symbols represent numerical results obtained from Eq. (11) using the population dynamics algorithm with $M=10^{6}$ (vertical bars indicate the standard deviation of the mean computed over 1010 independent runs). Solid lines are analytic predictions derived from the moments of the leading eigenvector [Eqs. (47) and (48)].
@@
of the infection rates or coupling strengths. Depending on the network parameters, Eqs. (47) and (48) diverge due to vanishing denominators Metz and Neri (2021). In particular, $\mathcal{I}^{(4)}(R_{1})$ diverges at $c=\mu_{4}^{\frac{1}{3}}$. This naturally raises the question of whether $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$ exhibit a similar behaviour near $c=\lambda^{-1}$.
In figure 5, we show $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$ as functions of $\sigma$ near the epidemic threshold. As a comparison, this figure also displays Eqs. (47) and (48) at $c=\lambda^{-1}$, demonstrating that $\mathcal{I}^{(2)}(R_{1})$ and $\mathcal{I}^{(4)}(R_{1})$ diverge at $\sigma_{*}=\sqrt{\lambda(1-\lambda)}$ and $\sigma_{4}\simeq 0.36$, respectively, for the $\Gamma$-distribution of infection rates. The results for $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$ are overall consistent with the analytic predictions obtained from the moments of $\bm{R}_{1}$. The discrepancies observed in figure 5 stem from the slow convergence of the population dynamics algorithm near the thresholds $\sigma_{*}$ and $\sigma_{4}$, which makes it difficult to accurately determine the distribution $\mathcal{P}(\rho)$ from the solutions of Eq. (11), as shown in appendix A. In addition, we note that $\mathcal{I}^{(4)}(\rho)$ remains bounded for $\sigma_{4}<\sigma<\sigma_{*}$, a consequence of the finite number of stochastic variables used to discretize $\mathcal{P}(\rho)$ in the population dynamics method.
@@
IV.4 Stochastic simulations
In this section, we compare our theoretical results in the limit $N\rightarrow\infty$, obtained within the QMF approximation, with numerical results from stochastic simulations of the SIS model Kiss et al. (2017); Cota and Ferreira (2017); Wu et al. (2019) on directed networks. The simulations are performed using the well-known Gillespie algorithm Kiss et al. (2017).
In the stochastic simulations of the SIS model, the dynamical state of each node $i$ is represented by a binary variable $s_{i}\in\{0,1\}$ ($i=1,\dots,N$). If node $i$ is infected, then $s_{i}=1$; otherwise, $s_{i}=0$. The uniform recovery rate $R_{i}$ of node $i$ is given by $R_{i}=s_{i}$, while the infection rate of vertex $i$ reads
$I_{i}=(1-s_{i})\sum_{j\in\partial_{i}^{\rm in}}\lambda_{ij}s_{j},$ (50)
where $\partial_{i}^{\rm in}$ denotes the in-neighbourhood of $i$. The coefficient $\lambda_{ij}>0$ is the directed coupling strength from node $j$ to $i$, representing the rate at which $j$ infects $i$. The variables $\{\lambda_{ij}\}_{i,j=1,\dots,N}$ are sampled from a distribution $P_{\lambda}(x)$ which, together with the degree distribution $p_{k,\ell}$, defines the network ensemble (see section II). Thus, infected nodes recover at a unit and constant rate, while susceptible nodes may become infected through contact with their in-neighbourhood.
The time increment between two consecutive updates in the simulation is sampled from an exponential distribution with rate equal to the total rate of change, $\sum_{i=1}^{N}(R_{i}+I_{i})$. To perform a single update in the system, we first select the type of event according to the total recovery and infection rates. If the selected event is a recovery, an infected node is chosen at random and its state is updated to susceptible. If the event is an infection, a susceptible node becomes infected with a probability proportional to its infection rate. To select which susceptible node becomes infected, we employ a rejection sampling method, i.e., we choose a node at random and accept the choice with probability $I_{i}/\sum_{j=1}^{N}I_{j}$. The prevalence at a given time is computed as
$\langle s\rangle=\frac{1}{N}\sum_{i=1}^{N}s_{i}.$ (51)
The stochastic dynamics of the SIS model on a finite network always converges to the absorbing state for sufficiently long times Sander et al. (2016); Cota and Ferreira (2017), which leads to computational difficulties in determining the epidemic threshold. To circumvent this issue, we employ the quasistationary method Sander et al. (2016), where the prevalence at time $t$ is computed by averaging Eq. (51) over the subset of realizations that have not reached the absorbing state.
Figure 7 compares theoretical results obtained from the solutions of Eq. (11) with stochastic simulations for different system sizes $N$. Finite size effects are small in the simulations, and the theoretical predictions for the prevalence, based on the QMF approximation, are in qualitative agreement with the simulation results. Since the QMF equations neglect pairwise dynamical correlations, the prevalence computed from Eq. (11) is systematically larger than the simulation results. Accordingly, the theoretical epidemic threshold is consistently smaller than the one observed in simulations. Nevertheless, figure 7 shows that the agreement between simulations and the QMF approximation improves as either $c$ increases or $\sigma$ decreases, i.e., in regimes where fluctuations in the contact network structure become less pronounced.
Figure 7: Comparison between theoretical results in the limit $N\rightarrow\infty$ (dashed lines) and stochastic simulation results (symbols) for the stationary prevalence of the SIS model on directed networks. The indegrees and outdegrees are independently drawn from a Poisson distribution with mean $c$, while the infection rates or coupling strengths $\lambda_{ij}$ follow a $\Gamma$-distribution with mean $\lambda=1/2$ and standard deviation $\sigma$. The theoretical results are obtained by solving Eq. (11) using the population dynamics algorithm with $M=10^{6}$ stochastic variables, while the simulation results are computed by averaging the prevalence over 1010 independent realizations after a long simulation time. The simulations are performed using the Gillespie algorithm (see the main text). Left panel (a): $\sigma=0.2$. Right panel (b): $c=5$.
Figure 8: Stochastic simulation results (symbols) for the prevalence of the SIS model on directed networks with two different degree distributions and a $\Gamma$-distribution of infection rates with mean $\lambda=1/2$ and standard deviation $\sigma$. The simulations are performed using the Gillespie algorithm (see the main text), and the results are obtained by averaging the prevalence over 110110 independent realizations of size $N=10^{4}$ after a long simulation time. The solid lines are a guide to the eye. Left panel (a): $\sigma=0$. Right panel (b): $c=3$.
In figure 8, we present simulation results for the prevalence across the epidemic threshold for $N=10^{4}$ and two different degree distributions. These results suggest that the epidemic threshold is indeed independent of the degree distribution, in qualitative agreement with the predictions derived from Eq. (11) (see figures 3 and 4). Thus, even though our theoretical findings are based on the QMF approximation, Eqs. (2), this simplified description appears to capture some essential aspects of the phase diagram.
@@ V Final remarks
In this work, we have determined the phase diagram of the SIS model on directed complex networks within the quenched mean-field approximation. By combining random-matrix results with an analytic approach for the distribution of stationary infection probabilities, we have computed the epidemic threshold as a function of the mean degree $c$ and the standard deviation $\sigma$ of the infection rates or coupling strengths defining the contact network. Our results show that the SIS model exhibits a transition between the absorbing and endemic phases provided $c\geq\lambda^{-1}$, where $\lambda$ is the average infection rate.
Remarkably, the critical line is independent of the degree distribution but it depends strongly on the distribution of infection rates. When $\sigma<\sigma_{*}=\sqrt{\lambda(1-\lambda)}$, both the epidemic threshold and the critical behaviour are governed by the leading eigenpair of the contact network Metz and Neri (2021) and are independent of the infection-rate distribution. In contrast, for $\sigma>\sigma_{*}$, the fluctuations of infection rates have a pronounced effect on the critical line. While the SIS model undergoes an absorbing phase transition for a $\Gamma$-distribution of infection rates as $\sigma$ increases, it remains in the endemic phase for a Pareto distribution. This striking difference is explained by the large fraction of near-zero infection rates generated by the $\Gamma$-distribution, which suppresses the endemic state for large $\sigma$. Together, these results provide a systematic characterization of how network heterogeneity shapes the phase diagram of the SIS model on directed networks in the quenched mean-field approximation.
@@
We have compared our theoretical results for $N\rightarrow\infty$ with stochastic simulations performed using the Gillespie algorithm Kiss et al. (2017); Cota and Ferreira (2017). The epidemic threshold computed from the solutions of Eq. (11) consistently lies below the value obtained from simulations. This quantitative discrepancy is not unexpected, since Eq. (11) is derived from the quenched mean-field (QMF) approximation, which neglects pairwise dynamical correlations Mata and Ferreira (2013). Nevertheless, the simulation results suggest that the exact epidemic threshold is independent of the degree distribution, in agreement with our theoretical predictions. Overall, this comparison with stochastic simulations indicates that the QMF approximation captures the main qualitative features of the phase diagram.
We remark that our results are strictly valid in the limit $N\rightarrow\infty$, as they rely on the numerical solutions of Eq. (11). Consequently, our approach does not address how the infection probabilities or the number of potentially infected nodes in the QMF theory scale with the system size, which could in principle clarify, for instance, why the dimensionless second moment remains finite while the IPR diverges. A detailed analysis of this issue and its relation with the epidemic survival time are interesting directions for future work. Furthermore, it would be interesting to unveil the mechanism sustaining the endemic state in directed graphs and compute the finite-size epidemic threshold by developing, for instance, a directed version of the scaling theory in Castellano and Pastor-Satorras (2020).
The SIS model on synthetic networks provides a theoretical framework to qualitatively investigate the effect of the heterogeneous structure of contact networks on epidemic spreading. However, we emphasize that, from a practical standpoint, this setting is highly idealized and insufficient for making quantitative predictions about real-world spreading phenomena. From a methodological perspective, we have introduced a general analytic framework to obtain an equation for the distribution of fixed-point states in the limit $N\rightarrow\infty$. The method is not restricted to the SIS model and can be extended to study the non-equilibrium fixed-points of other dynamical systems on directed networks, including the Kuramoto model Acebrón et al. (2005) and firing-rate models of neural networks Sompolinsky et al. (1988). Moreover, since the approach is based on the cavity method, it can also incorporate other important network features, such as short loops and degree-degree correlations, opening promising avenues of future research.
@@ Acknowledgements.
Appendix A Convergence of the population dynamics algorithm
In this appendix, we present numerical results for the influence of both the population size $M$ and the number of iterations $N_{\rm iter}$ in the population dynamics algorithm near the transition point $\sigma_{4}\simeq 0.36$ of figure 5. Figure 9 shows the difference between the analytic results for $N\rightarrow\infty$, Eqs. (47) and (48), and the numerical results for $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$ obtained from the solutions of Eq. (11). For relatively small number of iterations, there is a significant discrepancy between analytic and numerical results, which is not evident in figure 5 due to the logarithmic scale of the vertical axis. However, the data in figure 9 indicate that, despite the fluctuations, the numerical results consistently converge to the theoretical predictions as we increase both $M$ and $N_{\rm iter}$, giving further support to the analytic curves in figure 5. This critical slowing down of the population dynamics method close to $\sigma_{4}$ makes accurate calculations of higher moments of $\mathcal{P}(\rho)$ computationally very expensive.
Figure 9: Dimensionless second ($\mathcal{I}^{(2)}(\rho)$) and fourth ($\mathcal{I}^{(4)}(\rho)$) moments of the infection probabilities. The panels show the difference between the analytic results, $\mathcal{I}^{(2)}(R_{1})$ and $\mathcal{I}^{(4)}(R_{1})$ (see Eqs. (47) and (48)), and the numerical results for $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$, obtained from the solutions of Eq. (11) using the population dynamics algorithm with population size $M$ and total number $N_{\rm iter}$ of iterations (vertical bars indicate the standard deviation of the mean values computed over different samples). These results are for parameters near the critical point in figure 5: Poisson indegrees, $\Gamma$-distribution of infection rates with mean $\lambda=1/2$ and standard deviation $\sigma=0.3$, and $c-\lambda^{-1}=10^{-4}$.